### v1

# Attractor learning for spatiotemporally chaotic dynamical systems using echo state networks with transfer learning

Abstract
In this paper, we explore the predictive capabilities of echo state networks (ESNs) for the generalized Kuramoto-Sivashinsky (gKS) equation, an archetypal nonlinear PDE that exhibits spatiotemporal chaos. We introduce a novel methodology that integrates ESNs with transfer learning, aiming to enhance predictive performance across various parameter regimes of the gKS model. Our research focuses on predicting changes in long-term statistical patterns of the gKS model that result from varying the dispersion relation or the length of the spatial domain. We use transfer learning to adapt ESNs to different parameter settings and successfully capture changes in the underlying chaotic attractor.
Keywords: echo state networks, reservoir computing, generalized Kuramoto-Sivashinsky equation, transfer learning
1 Introduction
Over the past decade, there has been a substantial increase in research focused on the application of machine learning to dynamical systems. Neural networks have played a particularly important role, as a variety of promising techniques based on neural network architectures have been developed. In particular, reservoir computing [31, 58, 59] has been successfully utilized to learn and predict the dynamics of complex systems, including chaotic models. Reservoir computing refers to a broad class of recurrent neural networks, wherein the reservoir has its own nonlinear dynamics and only a linear output layer is trained to match the observational data.
In this paper, we explore the predictive capabilities of echo state networks (ESNs) in the context of spatiotemporal chaos. An echo state network is a simple reservoir computing architecture with the echo state property [28, 29, 39, 7, 30, 21, 58]. We introduce a novel methodology that integrates ESNs with transfer learning (TL). Our goal is to demonstrate that this novel methodology significantly enhances predictive performance when predicting changes in long-term statistical behavior that result from varying model parameters.
We describe our approach in the context of flows on metric spaces. Let $(M,d)$ be a metric space and let $\Phi:M\times[0,\infty)\to M$ be a continuous map that defines a flow on $M$. That is, we have $\Phi^{t+s}=\Phi^{t}\circ\Phi^{s}$ for all $s,t\geqslant 0$, where $\Phi^{t}(\cdot)=\Phi(\cdot,t)$. Think of $M$ as a function space and $\Phi$ as a flow on this function space generated by an evolution partial differential equation. Research at the interface of machine learning and dynamical systems has asked the following question. Can $\Phi$ be learned for predictive purposes?
The single-orbit approach asks the machine learning architecture to predict individual trajectories of $\Phi$ once training is complete. Echo state networks have been successfully used to predict trajectories of complex models, including chaotic systems [46, 19, 25, 22, 42, 11, 14, 23]. It is challenging, though, to prove rigorous mathematical lower bounds for the length of time the single-orbit predictions remain accurate. Alternatively, one can ask the machine learning architecture to learn and then predict statistical properties of the dynamics. This approach has also succeeded. Indeed, echo state networks can learn statistical properties of attractors of chaotic systems [46, 45, 38, 56, 37, 3, 48, 51].
In this paper, we focus on the following problem. Suppose that the flow $\Phi$ depends on a parameter $\gamma$. We write $\Phi_{\gamma}$ to indicate this dependence. For example, $\gamma$ could be a parameter in the PDE that generates the flow $\Phi_{\gamma}$. Suppose that for each $\gamma$ in some parameter set $\Gamma$, the flow $\Phi_{\gamma}$ admits an attractor that supports an ergodic invariant measure $\mu_{\gamma}$. Suppose further that for each $\gamma\in\Gamma$, $\mu_{\gamma}$ describes the asymptotic distribution of the orbit of a typical $x\in M$, meaning that
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\delta_{\Phi_{\gamma}^{s}(x)}=\mu_{\gamma}$ (1)
in the weak-$\ast$ topology. In this setting, we ask two questions. First, if we fix a parameter value $\gamma_{1}\in\Gamma$ and train an echo state network using trajectories from $\Phi_{\gamma_{1}}$, can the echo state network then predict statistical properties of $(f\circ\Phi_{\gamma_{1}}^{t})_{t\geqslant 0}$, where $f:M\to\mathbb{R}^{n}$ is a given observable? Second, assuming an affirmative answer to the first question, how might we predict the statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ with respect to a different parameter value $\gamma_{2}\in\Gamma$?
The point of this paper is to answer these two questions. Our approach is as follows. Suppose the echo state network has been trained using trajectories from $\Phi_{\gamma_{1}}$ and can successfully predict statistical properties of $(f\circ\Phi_{\gamma_{1}}^{t})_{t\geqslant 0}$. Since invariant measures of dynamical systems can depend sensitively on parameters, the invariant measure $\mu_{\gamma_{2}}$ associated with a different parameter value $\gamma_{2}$ might differ meaningfully from $\mu_{\gamma_{1}}$. Consequently, we would not expect the echo state network to successfully predict statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ when trained using trajectories from $\Phi_{\gamma_{1}}$. We conjecture that the predictive capability of the echo state network can be ‘transferred’ from the $\gamma_{1}$-system to the $\gamma_{2}$-system using transfer learning [27, 11]. That is, the echo state network will successfully predict statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ after we update its training using a small new dataset from the new regime.
We show that our proposed approach succeeds for the generalized Kuramoto-Sivashinsky (gKS) equation, an archetypal nonlinear PDE that exhibits spatiotemporal chaos. We train an echo state network to accurately predict the statistics of gKS dynamics. We then show that transfer learning allows the echo state network to retain predictive power as we vary a parameter in the dispersion relation of the gKS equation or the length of the physical domain.
Transfer learning is a popular concept in many areas of machine learning, including speech recognition and image processing. However, transfer learning for echo state networks has received comparatively little attention in the literature. We conclude that transfer learning can allow echo state networks to maintain predictive power as model parameters change. Such flexibility is important when statistical properties of the dynamics depend meaningfully on model parameters.
2 KS and gKS equations
2.1 Background
The one-dimensional Kuramoto-Sivashinsky (KS) equation [36, 49] has become a canonical model for the study of spatiotemporal chaos. The one-dimensional KS equation is given by
$\frac{\partial u}{\partial t}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}+u\frac{\partial u}{\partial x}=0,$ (2)
with the periodic boundary condition $u(x+L,t)=u(x,t)$ for all $x$ and every $t\geqslant 0$. The KS equation is well-behaved as an infinite-dimensional dynamical system. In particular, the KS model admits a finite-dimensional attractor and is therefore in some sense equivalent to a finite-dimensional dynamical system [13, 20, 43]. See [52] for an overview. Moreover, the KS model admits an inertial manifold that contains the global attractor [32, 12, 53]. The KS equation has been studied extensively, both analytically and computationally [35, 26, 50, 41].
The parameter $L$ (the length of the spatial domain) plays the role of a bifurcation parameter for the KS model. This parameter determines the number of linearly unstable Fourier modes, $S_{\mathrm{unst}}$. In addition, the maximum of the coefficient of the linearized equation occurs at $M_{\mathrm{unst}}$. We refer to $M_{\mathrm{unst}}$ as ”the most unstable wavenumber”. These quantities are given by
$S_{\mathrm{unst}}=\left\lfloor\frac{L}{2\pi}\right\rfloor,\qquad M_{\mathrm{unst}}=\frac{L}{2\sqrt{2}\pi}.$ (3)
[…]
The one-dimensional generalized Kuramoto-Sivashinsky equation is obtained from (2) by adding a parametric dispersion term (third derivative). The gKS equation is given by
$\frac{\partial u}{\partial t}+\frac{\partial^{2}u}{\partial x^{2}}+\gamma\frac{\partial^{3}u}{\partial x^{3}}+\frac{\partial^{4}u}{\partial x^{4}}+u\frac{\partial u}{\partial x}=0,$ (4)
where $\gamma$ is a parameter. The boundary conditions are again periodic.
The dynamics of the gKS equation have been studied extensively [34, 40, 33, 5, 8, 18, 15, 55, 16, 54, 47, 9]. The parameter $\gamma$ has a strong influence on gKS dynamics. For positive values of $\gamma$ near zero, the gKS model exhibits spatiotemporal chaos, like the KS model does. As $\gamma$ increases, the gKS model exhibits a transition toward less chaotic or even non-chaotic dynamics (see [24] for a detailed numerical study). In the limit $\gamma\to\infty$, the gKS equation is equivalent to the integrable Korteweg-de Vries (KdV) equation [55].
In this paper, we focus on statistical prediction as the parameters $L$ and $\gamma$ vary.
2.2 Numerical method
We use numerical solutions of the gKS equation to train ESNs. In this subsection, we describe the numerical method we use to integrate the gKS equation.
We use a uniform grid on $[0,L]$ and apply finite differences to discretize the derivatives in the gKS equation. In particular, we use an explicit method to discretize nonlinear fluxes and treat linear terms implicitly. We utilize Euler time-stepping. Therefore, the discrete version of the gKS equation reads
$\frac{u_{j}^{n+1}-u_{j}^{n}}{\delta t}=-\frac{F_{j+1/2}^{n}-F_{j-1/2}^{n}}{\Delta x}+f({\mathbf{u}}^{n+1}),$ (5)
where $F_{j+1/2}=\frac{1}{6}((u_{j}^{n})^{2}+u_{j}^{n}u_{j+1}^{n}+(u_{j+1}^{n})^{2})$, $f({\mathbf{u}}^{n+1})$ represents the discretization of the linear second, third, and fourth derivatives, and ${\mathbf{u}}^{n+1}=\{u_{j}^{n+1}:j=1,\ldots,N_{x}\}$ is the discrete solution vector at time $n+1$. The second, third, and fourth derivatives are discretized using standard $O(\Delta x^{2})$ central finite-difference formulas.
3 Echo state networks and transfer learning
Recurrent neural networks can be challenging to train. Reservoir computing has emerged as an elegant solution to the training challenge [28, 30, 39]. This is so because the reservoir has its own nonlinear dynamics and only a linear output layer is trained to match the observational data. Reservoir computers can successfully predict chaotic dynamics [31, 45, 10, 6, 42, 57]. They have been successfully applied in a variety of fields, including climate and weather prediction [42, 4], turbulent convection [44], and others [10, 57].
In this paper, we use the following standard ESN architecture. The ESN consists of an input layer, a reservoir of $D$ neurons, and an output layer. Connections between the neurons are encoded by an adjacency matrix $A$. The reservoir processes successive inputs, where the reservoir state at time $t+\Delta t$, $r(t+\Delta t)\in\mathbb{R}^{D}$, is a nonlinear function of the previous reservoir state, $r(t)\in\mathbb{R}^{D}$, and the driving input, $X(t)$, given by
$r(t+\Delta t)=f(Ar(t)+W_{\mathrm{in}}X(t)).$ (6)
Here $f=\tanh:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is the element-wise activation function. Let $N_{x}$ denote the dimension of the input. Matrices $A\in\mathbb{R}^{D\times D}$ and $W_{\mathrm{in}}\in\mathbb{R}^{N_{x}\times D}$ remain constant throughout the training process. During training, only the weights of the output layer, $W_{\mathrm{out}}\in\mathbb{R}^{D\times N_{x}}$, are tuned. Predictions are given by
$X(t+\Delta t)=W_{\mathrm{out}}\phi(r(t+\Delta t)),$ (7)
where $\phi$ is the nonlinear transformation [10, 45] given component-wise by
$\tilde{r}_{j}(t)=\phi(r_{j}(t))=\begin{cases}r_{j}^{2}(t)&\text{if $j$ is even}\\ r_{j}(t)&\text{if $j$ is odd}.\end{cases}$ (8)
We found in our experiments that this nonlinear transformation slightly improves the quality of predictions for individual trajectories.
The input matrix $W_{\mathrm{in}}$ is generated component-wise from a uniform distribution on $[-\beta_{1},\beta_{1}]$. The adjacency matrix is generated as $A=\beta_{2}W_{0}/\lambda$, where $W_{0}$ is a sparse matrix (typically turning on less than $10\%$ of the possible connections), $\lambda$ denotes the largest eigenvalue of $W_{0}$, and $\beta_{2}$ is a scaling parameter.
The weights of $W_{\mathrm{out}}\in\mathbb{R}^{{D}\times N_{x}}$ are determined through a linear regression with $L^{2}$-regularization by
$W_{\mathrm{out}}=\operatorname*{arg\,min}_{W}\|W\mathbf{R}-\mathbf{X}\|_{2}^{2}+\mu\|W\|_{2}^{2},$ (9)
where $\mathbf{X}$ is the training data (trajectories), and $\mathbf{R}$ denotes the reservoir timeseries (with $A$ and $W_{\mathrm{in}}$ fixed). $W_{\mathrm{out}}$ can be expressed explicitly as
$W_{\mathrm{out}}=\mathbf{X}\mathbf{R}^{\prime}(\mathbf{R}\mathbf{R}^{\prime}+\mu I)^{-1}.$ (10)
Here, $\mathbf{R}^{\prime}$ denotes the transpose of $\mathbf{R}$. Throughout this paper, we use ESN parameters $D=5000$, $\beta_{1}=0.01$, $\beta_{2}=0.1$, and $\mu=0.000005$, unless stated otherwise.
Transfer learning [27, 11] is a machine-learning approach wherein knowledge in a source domain is used to improve predictive performance in a target domain. In our context, the source and target domains are simply two sets of gKS parameters. Our central premise for linking transfer learning to ESNs is the following. Once the ESN is well trained in the source domain, it will retain predictive power in the target domain after receiving a training update using a small transfer-learning dataset from the target domain.
Writing $\mathbf{X}_{TL}$ for the transfer-learning dataset, the update to the output layer is obtained by solving the optimization problem given by
$\delta W=\operatorname*{arg\,min}_{\delta W}\left\|(W_{\mathrm{out}}+\delta W)\mathbf{R}_{TL}-\mathbf{X}_{TL}\right\|_{2}^{2}+\alpha\|\delta W\|_{2}^{2}.$ (11)
Here, $\alpha$ is the rate of transfer learning and $\mathbf{R}_{TL}$ is the (time-dependent) state of the reservoir (with the same number of time-snapshots as $\mathbf{X}_{TL}$). The update to the output layer and new output layer are given explicitly by
$\delta W$ $=\left(\mathbf{X}_{TL}\mathbf{R}_{TL}^{\prime}-W_{\mathrm{out}}\mathbf{R}_{TL}\mathbf{R}^{\prime}_{TL}\right)\left(\mathbf{R}_{TL}\mathbf{R}^{\prime}_{TL}+\alpha I\right)^{-1},$ (12)
$W_{\mathrm{out}}^{\mathrm{new}}$ $=W_{\mathrm{out}}+\delta W.$
Here, $W_{\mathrm{out}}$ and $W_{\mathrm{out}}^{\mathrm{new}}$ are the output layer in the source and target domains, respectively.
Transfer learning has been used with ESNs to successfully predict individual trajectories for shallow-water equations [11, 1] and the gKS equation [1]. In this paper, we show that transfer learning allows us to successfully track changes in the statistical signature of the gKS equation as either the size of the spatial domain or the dispersion relation vary.
4 Results
We focus on an important statistical feature of gKS dynamics, namely the power spectrum. The power spectrum is defined as the map from $\mathbb{Z}$ into $\mathbb{R}$ given by
$k\mapsto e_{k}=\frac{1}{T}\int\limits_{0}^{T}|{\hat{u}}_{k}(t)|^{2}\,\mathrm{d}t,$ (13)
where $\hat{u}_{k}(t)$ is coefficient $k$ at time $t$ of the Fourier-series representation of the solution to the gKS equation,
$u(x,t)=\sum_{k\in\mathbb{Z}}\hat{u}_{k}(t)e^{(2\pi ik/L)x}.$ (14)
The quantity $e_{k}$ represents the average energy in Fourier wavenumber $k$. Since the solution $u(x,t)$ of the gKS equation is real-valued, the Fourier coefficients satisfy ${\hat{u}}_{-k}={\hat{u}}_{k}^{*}$. We therefore consider only nonnegative wavenumbers.
Throughout Section 4, we work with the gKS equation in the chaotic regime ($\gamma=0$ or $\gamma$ is small and positive). We assume the gKS equation admits an attractor that supports an ergodic measure. We assume that this ergodic measure describes the distribution of the orbit of the initial data we select. (As $\gamma$ increases, the gKS equation eventually becomes less chaotic [24]. In particular, the power spectrum depends on the initial data when $\gamma$ is sufficiently large.)
We first consider the KS equation and we vary the parameter $L$. Varying $L$ is equivalent to varying the number of linearly unstable Fourier modes. We show that for each of several values of $L$, the ESN accurately predicts the power spectrum of the KS equation with spatial domain $[0,L]$, once trained using KS trajectories from the same spatial domain. We then show that transfer learning is effective for various tuples $(L_{1},L_{2})$ of length parameters. That is, the ESN accurately predicts the power spectrum of the KS equation with parameter $L_{2}$ after a two-step training process. First, the ESN is trained using trajectories of the KS equation with parameter $L_{1}$. Second, the training is updated using a small amount of trajectory information from the KS equation with parameter $L_{2}$.
We conclude Section 4 by carrying out the program in the previous paragraph for the gKS equation. This time, we fix $L$ and vary the parameter $\gamma$ in the dispersion relation. We show that when $\gamma$ is small, the ESN accurately predicts power spectra and transfer learning is effective.
4.1 Resolution study
In [1, Section 5.1.3], Alam analyzed numerically how spatial resolution affects the performance of ESNs when using them to predict the power spectrum of the gKS equation. In particular, Alam considered $L=43$ and $\gamma=0.1$ and analyzed how numerically computed power spectra depend on the spatial resolution. In addition, Alam investigated how averaging time affects the accuracy of the computation of power spectra. Alam determined that spatial resolution $N_{x}=256$, averaging time $T=1000$ Lyapunov times, integration timestep $\delta t=0.001$, and snapshot sampling timestep $\Delta t=0.25$ are sufficient to accurately reproduce the power spectrum of the gKS equation in the chaotic regime. We therefore use these parameter values for the simulations in this paper, unless stated otherwise.
4.2 Predicting the power spectrum of the KS equation
4.2.1 Predicting the power spectrum of the KS equation without transfer learning
We show that for each of several values of $L$, the ESN accurately predicts the power spectrum of the KS equation with spatial domain $[0,L]$, once trained using KS trajectories from the same spatial domain. We consider four values of $L$ that correspond to four different numbers of linearly unstable Fourier modes. Table 1 lists each choice of $L$, together with the corresponding number of linearly unstable Fourier modes ($S_{\mathrm{unst}}$), the most unstable wavenumber ($M_{\mathrm{unst}}$), and an approximation of $\lambda_{\max}$, the maximal Lyapunov exponent. We use the approximation for the Lyapunov exponents ($\lambda_{0}\geqslant\lambda_{1}\geqslant\cdots\geqslant\lambda_{i}\geqslant\cdots$) given in [17] by
$\lambda_{i}\left(L\right)\approx 0.093-0.94\left(i-0.39\right)/L.$ (15)
Table 1: Values of $L$ for which we predict the power spectrum of the KS equation with spatial domain $[0,L]$. For each value of $L$, we list the number of linearly unstable Fourier modes ($S_{\mathrm{unst}}$) given by (3), the most unstable wavenumber ($M_{\mathrm{unst}}$) given by (3), and an approximation of the maximal Lyapunov exponent (given in [17] and by (15)).
L $S_{\mathrm{unst}}$ $M_{\mathrm{unst}}$ $\lambda_{\max}$
22 3 2.48 0.1097
29 4 3.26 0.1056
35 5 3.94 0.1035
43 6 4.84 0.1015
We performed the following experiment for each value of $L$. We began with an untrained ESN. We generated $30$ KS trajectories of length $T=1000$ Lyapunov times using random initial data of the form
$u(x,0)=c_{1}\cos\left(\frac{p_{1}x\pi}{L}\right)+c_{2}\cos\left(\frac{p_{2}x\pi}{L}\right),$ (16)
where we independently selected $c_{1},c_{2}\sim\mathrm{Uniform}([0,1])$ and $p_{1},p_{2}\sim\mathrm{Uniform}(\{1,2,3,4,5,6\})$. We used $20$ of these trajectories to train the ESN and the other $10$ for prediction.
Before considering prediction of the power spectrum, it is natural to ask how well the trained ESN predicts individual KS trajectories. Alam [1, Section 4.3.2] investigated this question. He found that ESN predictions of individual trajectories remain accurate up to approximately $7$ Lyapunov times. This timescale is consistent with results published earlier [45]. Alam [1, Section 4.3.2] found that the timescale for which predictions of individual trajectories remains accurate decays approximately linearly with $L$. As $L$ increases, one can potentially deal with this decay by training a larger ESN or by training a coupled network of ESNs. Nevertheless, we are unaware of rigorous mathematical guarantees associated with single-orbit prediction of spatiotemporally chaotic dynamics using ESNs. That is, we are unaware of rigorous lower bounds on the number of Lyapunov times for which the single-orbit prediction remains accurate. This motivates our focus on the power spectrum. The idea is that once the ESN can predict the power spectrum, it can do so for a very long time. In particular, it can continue to accurately predict the power spectrum long after single-orbit predictions have become inaccurate.
Figure 1 demonstrates that for each of the four values of $L$ in Table 1, the ESN accurately predicts the power spectrum of the KS equation (once trained). Each panel compares the power spectrum obtained by direct numerical simulation (DNS) to the ESN prediction. We average over $10$ trajectories used for prediction to obtain each DNS-based curve.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Once trained on KS trajectories from spatial domain $[0,L]$, the ESN then accurately predicts the power spectrum of the KS equation with the same spatial domain. Each blue curve shows the power spectrum obtained from direct numerical simulation, averaged over $10$ trajectories used for prediction. Red curves show the ESN predictions of the power spectrum. Importantly, the ESN predictions of the power spectrum remain accurate long after ESN predictions of individual trajectories become inaccurate. Spatial domain sizes $L=22$ (top left), $L=29$ (top right), $L=35$ (bottom left), and $L=43$ (bottom right) correspond to four different numbers of linearly unstable Fourier modes (see Table 1). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.2.2 Predicting the power spectrum of the KS equation using transfer learning
When $L$ varies, it is computationally expensive to train a de novo ESN to predict the power spectrum for each new value of $L$. We therefore investigate the efficacy of transfer learning with respect to $L$ for the prediction of the power spectrum.
We consider $L=22$ and $L=43$, values of $L$ that correspond to the presence of $3$ and $6$ linearly unstable Fourier modes, respectively. Importantly, the power spectrum for $L=43$ differs considerably from the power spectrum for $L=22$ (Figure 2, blue and red curves, respectively).
We performed the following experiment. First, we trained the ESN in the $L=22$ regime using $20$ trajectories from that regime. Second, we set the transfer learning rate to $\alpha=0.005$ and then updated the training of the ESN using trajectory data from the $L=43$ regime. (We tested several values of $\alpha$, including $0.1$, $0.01$, $0.005$, $0.002$, and $0.001$. Empirically, $\alpha=0.005$ worked well.) We report the amount of TL data used for the training update as a percentage of the $20$ trajectories from the $L=43$ regime used for training.
In the absence of transfer learning, the ESN (trained only on trajectories from the $L=22$ regime) performs poorly in the $L=43$ regime for both single-orbit prediction and the prediction of the power spectrum [1]. However, transfer learning dramatically improves the prediction of the power spectrum in the $L=43$ regime. Figure 2 illustrates this dramatic improvement as a function of amount of TL data. The blue curve (same curve in each panel) is the power spectrum in the $L=43$ regime, computed by averaging over the $10$ DNS trajectories from this regime used for prediction. This curve serves as ground truth. The red curve (same curve in each panel) is the power spectrum for $L=22$ predicted by the ESN after having been trained on $20$ DNS trajectories from the $L=22$ regime. Comparing the red and blue curves, we see that the power spectrum for $L=22$ differs considerably from that for $L=43$. The green curves illustrate the ESN prediction of the power spectrum for the $L=43$ regime after transfer learning of level $10\%$ (left), $25\%$ (center), and $50\%$ (right). The quality of the prediction based on transfer learning rapidly improves as the level of TL data increases. In particular, the prediction is nearly perfect at level $50\%$ (right).
Figure 2: Transfer learning allows the ESN to accurately predict the power spectrum of the KS equation as the size of the spatial domain varies. Curves illustrate power spectra. DNS-L43: Power spectrum obtained by averaging over the $10$ trajectories in the $L=43$ regime used for prediction. ESN-L22: Power spectrum for the $L=22$ regime predicted by the ESN, once it is trained on $20$ trajectories from the $L=22$ regime. ESN-TL: Power spectrum for the $L=43$ regime predicted by the ESN after transfer learning of level $10\%$ (left), $25\%$ (middle), and $50\%$ (right). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.3 Predicting the power spectrum of the gKS equation
We consider the gKS equation when the parameter $\gamma$ in the dispersion relation is small. We assume that for $\gamma$ small, the gKS dynamics admit an attractor that supports an ergodic invariant measure. We assume that this invariant measure describes the asymptotic distribution of all of the initial data that we select. We show that for $\gamma$ small and fixed, the ESN accurately predicts the power spectrum of the gKS equation for this fixed value of $\gamma$, once trained on trajectories generated using this fixed value of $\gamma$. We then show that transfer learning is remarkably effective for small $\gamma$. Throughout Section 4.3, we set $L=43$.
4.3.1 Predicting the power spectrum of the gKS equation without transfer learning
We performed the following experiment for each of two values of $\gamma$, $\gamma=0$ and $\gamma=0.1$. We began with an untrained ESN. We generated $30$ gKS trajectories, each of length $T=1000$ Lyapunov times, via direct numerical simulation. We used $20$ of these to train the ESN and the other $10$ for prediction.
Figure 3 illustrates that for $\gamma=0$ (left) and $\gamma=0.1$ (right), the trained ESN accurately predicts the power spectrum of the gKS equation. Each red curve shows the ESN prediction of the power spectrum. Each blue curve shows the power spectrum obtained by averaging over the $10$ trajectories used for prediction. Importantly, the power spectrum of the gKS equation depends sensitively on $\gamma$ in the chaotic (small $\gamma$) regime. This observation serves as serious motivation for the use of transfer learning in this regime.
Refer to caption Refer to caption
Figure 3: Once trained, the ESN accurately predicts the power spectrum of the gKS equation. Left: $\gamma=0$. Right: $\gamma=0.1$. Each blue curve shows the power spectrum obtained by averaging over the $10$ DNS trajectories used for prediction. Each red curve shows the power spectrum predicted by the ESN, after the untrained ESN has been trained using $20$ trajectories. Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.3.2 Predicting the power spectrum of the gKS equation with transfer learning
Figure 3 demonstrates that the statistical behavior of the gKS equation depends sensitively on $\gamma$ in the small-$\gamma$ regime. We therefore ask if transfer learning can capture substantial changes in statistical behavior as $\gamma$ varies. To address this question, we performed the following experiment. We started with an untrained ESN and first trained it using $20$ trajectories from the $\gamma=0$ regime. We then updated the training using a small amount of TL data from the $\gamma=0.1$ regime at transfer learning rate $\alpha=0.005$.
Refer to caption Refer to caption
Figure 4: Transfer learning allows the ESN to accurately predict the power spectrum of the gKS equation as the dispersion parameter $\gamma$ varies. Blue curve (same curve in both panels): Power spectrum of the gKS equation in the $\gamma=0.1$ regime obtained by averaging over the $10$ DNS trajectories used for prediction. Green curve (same curve in both panels): Power spectrum predicted by the ESN after a first round of training using $20$ trajectories from the $\gamma=0$ regime and then a training update using TL data from the $\gamma=0.1$ regime at level $10\%$. Yellow curve: Power spectrum predicted by the ESN after training only on $20$ trajectories from the $\gamma=0$ regime. Red curve: Power spectrum predicted by an ESN that has been trained only on the TL data, namely $2$ trajectories from the $\gamma=0.1$ regime. Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
The result of this experiment is shown in Figure 4 (left). The blue curve shows the ground truth, namely the power spectrum in the $\gamma=0.1$ regime obtained by averaging over the $10$ trajectories from the $\gamma=0.1$ regime used for prediction. Using a transfer learning level of only $10\%$ ($2$ trajectories) produces an accurate ESN prediction of the power spectrum (green curve). For reference, the yellow curve shows the ESN-based prediction of the power spectrum when the ESN is trained only on $20$ trajectories from the $\gamma=0$ regime. Importantly, the accurate TL-based prediction of the power spectrum is valid over long timescales. By contrast, single-orbit predictions in the $\gamma=0.1$ regime based on transfer learning are valid for shorter timescales [1, Section 5.1.1].
As a control, we compare the prediction using transfer learning (Figure 4, green curve, ESN-TL) to a prediction from an ESN that has been trained only on the TL data, namely $2$ trajectories from the $\gamma=0.1$ regime (Figure 4, red curve, ESN*). We see that the ESN* prediction is inferior to the ESN-TL prediction in the eye metric. Discrepancies between these two predictions occur at low wavenumbers that contain most of the energy. Quantitatively, the curves shown in Figure 4 (right) contain total energies $2.92$ (blue, DNS), $2.80$ (red, ESN*), and $2.97$ (green, ESN-TL). This corresponds to relative error $4.1\%$ and $1.7\%$ for the ESN* and ESN-TL predictions, respectively. In the relative-error metric, the ESN-TL prediction is more than twice as accurate as the ESN* control.
5 Discussion
The interface of machine learning and dynamical systems is a rapidly developing field. Echo state networks have emerged as a powerful tool. For spatiotemporally chaotic systems, these networks are capable of both single-orbit and statistical prediction. In this paper, we have focused on predicting the long-term statistical behavior of an archetypal model, the gKS equation in the spatiotemporally chaotic regime.
We have shown that when the domain size $L$ and the dispersion parameter $\gamma$ are fixed, our ESN can learn and then accurately predict the power spectrum of the gKS equation. Further, we have shown that transfer learning allows one to efficiently and accurately track changes in the power spectrum of the gKS equation as $L$ or $\gamma$ varies. Importantly, our ESN did not blow up nor collapse during our long simulations. This is a great sign for the utility of ESNs with respect to long-term statistical prediction.
Looking to the future, ESNs can potentially be used as an inexpensive climate simulator or a chaotic heat bath. Importantly, ESNs allow relatively large timesteps when generating new trajectories, making them more efficient than traditional numerical integrators. Thus far, transfer learning for ESNs has received limited attention in the literature. We believe transfer learning is an important concept that can be applied to complex problems and allow ESNs to track parametric changes in spatiotemporally chaotic dynamics. In general, developing efficient models that can track parametric changes in dynamical systems is an important task. The methodology described in this paper aims at this goal.
6 Code
Our code for simulating the gKS equation and ESNs is located on GitHub [2].
7 Author contributions
All authors wrote the paper. Study design: WO, IT. Initial code development: IT. Code development, simulation, and visualization: MSA.